# An Information Criterion for Controlled Disentanglement of Multimodal Data

### v1

@@ Abstract
Multimodal representation learning seeks to relate and decompose information inherent in multiple modalities. By disentangling modality-specific information from information that is shared across modalities, we can improve interpretability and robustness and enable downstream tasks such as the generation of counterfactual outcomes. Separating the two types of information is challenging since they are often deeply entangled in many real-world applications. We propose Disentangled Self-Supervised Learning (DisentangledSSL), a novel self-supervised approach for learning disentangled representations. We present a comprehensive analysis of the optimality of each disentangled representation, particularly focusing on the scenario not covered in prior work where the so-called Minimum Necessary Information (MNI) point is not attainable. We demonstrate that DisentangledSSL successfully learns shared and modality-specific features on multiple synthetic and real-world datasets and consistently outperforms baselines on various downstream tasks, including prediction tasks for vision-language data, as well as molecule-phenotype retrieval tasks for biological data.
@@ 1 Introduction
Humans understand and interact with the world using multiple senses, each providing unique and complementary information essential for forming a comprehensive mental representation of the environment. To emulate human-like perception, multimodal representation learning (Bengio et al., 2013; Liang et al., 2022a) seeks to decipher complex systems by combining information from multiple modalities into holistic representations, showcasing significant applications across fields from vision-language (Yuan et al., 2021; Lu et al., 2019; Radford et al., 2021) to biology (Yang et al., 2021; Zhang et al., 2022; Wang et al., 2023a). Large multimodal representation learning models such as CLIP (Radford et al., 2021), trained through self-supervised learning, maximally capture the mutual information shared across multiple modalities, exploiting the assumption of multi-view redundancy (Tosh et al., 2021; Sridharan & Kakade, 2008). This property indicates that shared information between modalities is exactly what is relevant for downstream tasks.
However, the modality gap, rooted in the inherent differences in representational nature and information content across modalities (Liang et al., 2022b; Ramasinghe et al., 2024; Huh et al., 2024), leads to the misalignment between modalities and restricts the application of these methods in many real-world multimodal scenarios. This highlights the need for a disentangled representation space that captures both shared and modality-specific information effectively, elaborated as follows:
@@
Disentangled representation learning in multimodal data traces back to seminal works in Variational Autoencoders and Generative Adversarial Networks (Lee & Pavlovic, 2021; Daunhawer et al., 2021; Denton et al., 2017; Gonzalez-Garcia et al., 2018), aiming to isolate the underlying factors of variation within the data. Recent studies have increasingly explored self-supervised learning methods to capture both shared and modality-specific information (Liu et al., 2024b; Li et al., 2024; Zhang et al., 2024). In particular, information-theoretic approaches model modality-specific information as the complement of shared information, learning these representations either sequentially (Sanchez et al., 2020) or jointly (Liang et al., 2024; Pan et al., 2021). However, none of these methods provide rigorous definitions or theoretical guarantees on the optimality of disentangled representations, particularly under the multimodal scenario where the Minimum Necessary Information (MNI) point between the two modalities, as introduced in Fischer (2020), is often unattainable. In simple terms, the MNI criterion characterizes $Z$ as a representation of $X^{1}$ that perfectly captures all the important information needed to understand $X^{2}$, without containing any unnecessary details. Essentially, knowing $Z$ gives you the same understanding of $X^{2}$ as directly knowing $X^{1}$, and vice versa [^1] Formally, $I(Z;X^{1})=I(Z;X^{2})=I(X^{1};X^{2})$ (Fischer, 2020), which we further explain in Section 2.2.1.. However, the shared and modality-specific information is deeply entangled in a wide range of real-world applications, i.e. the shared and modality-specific components are intertwined and result in similar observations, leading to unattainable MNI. Figure 1 illustrates an example in the context of high-content drug screens. Specifically, the two modalities capture distinct but related aspects of drug mechanism and cancer cell viability, and extracting precise shared features from each modality is not feasible. We provide additional examples of unattainable MNI in Appendix A.
@@ 2.2.2 Optimal Shared Representations: MNI Attainable or not
We propose a definition of optimality for the shared representations that applies to both scenarios – when MNI is attainable or not, as defined in Equation 1:
@@ 2.2.3 Optimal Specific Representations: Ensuring Coverage and Disentanglement
Optimal modality-specific representations, $\hat{Z}_{s}^{1}$ and $\hat{Z}_{s}^{2}$, should capture information unique to each modality, being highly informative while minimizing redundancy with the shared representations. We hence define them via the optimization problems
@@ 2.3 DisentangledSSL: a Step-by-Step Optimization Algorithm
To achieve the optimal representations discussed in Section 2.2, we introduce a two-step training procedure. The first step focuses on optimizing the shared latent representation, ensuring it captures the minimum necessary information as close as possible. Building upon this, the second step utilizes the learned shared representations in step 1 to facilitate the learning of modality-specific representations. This sequential approach is formalized in the optimization objectives given in Equations 3 and 4:
@@ 2.3.1 Optimality Guarantee for the Learned Shared Representations
$L_{c}^{1}$ seeks representation $Z^{1}$ that maximizes the information shared between the modalities, i.e. $I(Z^{1};X^{2})$, while minimizing the information unique to each modality, i.e. $I(Z^{1};X^{1}|X^{2})$, to capture only the essential shared content. This aligns with the conditional entropy bottleneck (CEB) objective (Fischer, 2020), an extension of the IB Lagrangian $L=I(Z^{1};X^{2})-\tilde{\beta}I(Z^{1};X^{1})$ (see details in Appendix B). While it serves as a robust objective for learning shared information, its optimality remains underexplored in Fischer (2020), particularly when MNI is unattainable. Additionally, Fischer (2020) focuses on the supervised scenario where $X^{2}$ is the label of $X^{1}$, whereas we address the multimodal case with $X^{1}$ and $X^{2}$ being two data modalities, demonstrating its effectiveness in both attainable and unattainable MNI scenarios.
@@ 3 Experiments
We compare DisentangledSSL against a disentangled variational autoencoder baseline, DMVAE (Lee & Pavlovic, 2021), as well as various multimodal contrastive learning methods, including CLIP (Radford et al., 2021), which aligns different modalities to learn the shared representations, and FOCAL (Liu et al., 2024b) and FactorCL (Liang et al., 2024), which capture both shared and modality-specific information. Additionally, we evaluate a joint optimization variant of DisentangledSSL, JointOpt (Pan et al., 2021), to highlight the advantages of our step-by-step approach.
@@ 3.1 Simulation Study
Results. We assess the performance of the learned shared and modality-specific representations for different values of $\beta$ and $\lambda$, as shown in Figure 4. For comparison, we also evaluate JointOpt, DMVAE, and FOCAL across different hyperparameter settings. Specifically, for JointOpt, we vary both $a$ and $\lambda$, where $a$ controls the joint mutual information terms $I(\hat{Z}_{s}^{1},\hat{Z}_{c}^{2};X^{1})$ and $I(\hat{Z}_{s}^{2},\hat{Z}_{c}^{1};X^{2})$, and $\lambda$ adjusts the mutual information term between representations, similar to DisentangledSSL.
@@ 3.2 MultiBench
Results. We evaluate the linear probing accuracy of representations learned by each pretrained model, as shown in Table 1. FactorCL-emb refers to the embeddings of FactorCL before projection heads, while FactorCL-proj uses the concatenation of all projection head outputs. All models use representations (or concatenation of them, if applicable) with the same dimensionality. We present results for the combined shared and specific representations for FOCAL and JointOpt in Table 1, with results for individual representations in Appendix H.2. DisentangledSSL consistently outperforms baselines across datasets, demonstrating its ability to capture valuable information for downstream tasks. Combining shared and specific representations of DisentangledSSL improves performance in most cases, showing that both contribute to label prediction, while in MOSI and MUSTARD, only shared or specific information contributes significantly.
@@ 4 Related Work
Multimodal disentanglement in self-supervised learning. Recent studies have explored disentangled representation learning in multimodal self-supervised contexts. Liang et al. (2024) factorizes and optimizes mutual information bounds to capture the union of shared and unique features, while not penalizing for redundancy between latents. Zhang et al. (2024) learns disentangled representations of cell state for multimodal data in biological contexts. Liu et al. (2024b) and Li et al. (2024) use mutual information optimization with orthogonal/MMD regularizations for disentangling multimodal time-series signals and T-cell receptor segments, respectively. However, these methods lack a general framework for understanding information content, particularly when data modalities are deeply entangled in real-world applications.
@@ Acknowledgement
Xinyi Zhang, Sana Tonekaboni and Caroline Uhler were partially supported by the Eric and Wendy Schmidt Center at the Broad Institute, NCCIH/NIH (1DP2AT012345), ONR (N00014-22-1-2116) and DOE (DE-SC0023187). SG and SJ acknowledge the support of the NSF Award CCF-2112665 (TILOS AI Institute) and an Alexander von Humboldt fellowship. CW and TJ acknowledge support from the Machine Learning for Pharmaceutical Discovery and Synthesis (MLPDS) consortium, the DTRA Discovery of Medical Countermeasures Against New and Emerging (DOMANE) threats program, and the NSF Expeditions grant (award 1918839) Understanding the World Through Code.
@@ Appendix B Background on Information Bottleneck
$\mathcal{L}=I(Z^{1};X^{1})-\beta I(Z^{1};Z^{2})$ (6)

### v2

@@ Abstract
Multimodal representation learning seeks to relate and decompose information inherent in multiple modalities. By disentangling modality-specific information from information that is shared across modalities, we can improve interpretability and robustness and enable downstream tasks such as the generation of counterfactual outcomes. Separating the two types of information is challenging since they are often deeply entangled in many real-world applications. We propose Disentangled Self-Supervised Learning (DisentangledSSL), a novel self-supervised approach for learning disentangled representations. We present a comprehensive analysis of the optimality of each disentangled representation, particularly focusing on the scenario not covered in prior work where the so-called Minimum Necessary Information (MNI) point is not attainable. We demonstrate that DisentangledSSL successfully learns shared and modality-specific features on multiple synthetic and real-world datasets and consistently outperforms baselines on various downstream tasks, including prediction tasks for vision-language data, as well as molecule-phenotype retrieval tasks for biological data. The code is available at https://github.com/uhlerlab/DisentangledSSL.
@@ 1 Introduction
Humans understand and interact with the world using multiple senses, each providing unique and complementary information essential for forming a comprehensive mental representation of the environment. To emulate human-like perception, multimodal representation learning (Bengio et al., 2013; Liang et al., 2022a) seeks to decipher complex systems by combining information from multiple modalities into holistic representations, showcasing significant applications across fields from vision-language (Yuan et al., 2021; Lu et al., 2019; Radford et al., 2021) to biology (Yang et al., 2021; Zhang et al., 2022; Wang et al., 2023a). Large multimodal representation learning models such as CLIP (Radford et al., 2021), trained through self-supervised learning, maximally capture the mutual information shared across multiple modalities. These models exploit the assumption of multi-view redundancy (Tosh et al., 2021; Sridharan & Kakade, 2008), which indicates that shared information between modalities is exactly what is relevant for downstream tasks.
However, the misalignment between modalities restricts the application of these methods in real-world multimodal scenarios. Among various contributing factors, the modality gap, rooted in the inherent differences in representational nature and information content across modalities (Liang et al., 2022b; Ramasinghe et al., 2024; Huh et al., 2024), plays a significant role. This highlights the need for a disentangled representation space that captures both shared and modality-specific information:
@@
Disentangled representation learning in multimodal data traces back to seminal works in Variational Autoencoders and Generative Adversarial Networks (Lee & Pavlovic, 2021; Daunhawer et al., 2021; Denton et al., 2017; Gonzalez-Garcia et al., 2018), aiming to isolate the underlying factors of variation within the data. Recent studies have increasingly explored self-supervised learning methods to capture both shared and modality-specific information (Liu et al., 2024b; Li et al., 2024; Zhang et al., 2024). In particular, information-theoretic approaches model modality-specific information as the complement of shared information, learning these representations either sequentially (Sanchez et al., 2020) or jointly (Liang et al., 2024; Pan et al., 2021). However, none of these methods provide rigorous definitions or theoretical guarantees on the optimality of disentangled representations, particularly under the multimodal scenario where the Minimum Necessary Information (MNI) point between the two modalities, as introduced in Fischer (2020), is often unattainable. In simple terms, the MNI criterion characterizes $Z$ as a representation of $X^{1}$ that perfectly captures all the important information needed to understand $X^{2}$, without containing any unnecessary details. Essentially, knowing $Z$ gives you the same understanding of $X^{2}$ as directly knowing $X^{1}$, and vice versa [^1] Formally, $I(Z;X^{1})=I(Z;X^{2})=I(X^{1};X^{2})$ (Fischer, 2020), which we further explain in Section 2.2.1.. However, the shared and modality-specific information is deeply entangled in a wide range of real-world applications, i.e., the shared and modality-specific components are intertwined and result in similar observations, leading to unattainable MNI. Figure 1 illustrates an example in the context of high-content drug screens. Specifically, the two modalities capture distinct but related aspects of drug mechanism and cancer cell viability, and extracting precise shared features from each modality is not feasible. We provide additional examples of unattainable MNI in Appendix A.
@@ 2.2.2 Optimal Shared Representations: MNI Attainable or not
[…]
@@ 2.2.3 Optimal Specific Representations: Ensuring Coverage and Disentanglement
Optimal modality-specific representations, $\hat{Z}_{s}^{1}$ and $\hat{Z}_{s}^{2}$, should capture information unique to each modality, being highly informative while minimizing redundancy with the shared representations. We hence define them via the optimization problems:
@@ 2.3 DisentangledSSL: a Step-by-Step Optimization Algorithm
To achieve the optimal representations discussed in Section 2.2, we introduce a two-step training procedure. The first step focuses on optimizing the shared latent representation, ensuring it captures the minimum necessary information as close as possible. Building upon this, the second step utilizes the learned shared representations in step 1 to facilitate the learning of modality-specific representations. This sequential approach is formalized in the optimization objectives given in Equations 3 and 4, with the pseudocode provided in Appendix I:
@@ 2.3.1 Optimality Guarantee for the Learned Shared Representations
$L_{c}^{1}$ seeks representation $Z^{1}$ that maximizes the information shared between the modalities, i.e. $I(Z^{1};X^{2})$, while minimizing the information unique to each modality, i.e. $I(Z^{1};X^{1}|X^{2})$, to capture only the essential shared content. This aligns with the conditional entropy bottleneck (CEB) objective (Fischer, 2020) [^5] Note that $L_{c}^{1}$ is the Lagrangian formulation of the constraint optimization in Equation 1, among which the term $I(X^{1};X^{2})$ is irrelevant to the optimization target $Z_{1}$ and omitted from the objective., an extension of the IB Lagrangian $L=I(Z^{1};X^{2})-\tilde{\beta}I(Z^{1};X^{1})$ (see details in Appendix B). While it serves as a robust objective for learning shared information, its optimality remains underexplored in Fischer (2020), particularly when MNI is unattainable. Additionally, Fischer (2020) focuses on the supervised scenario where $X^{2}$ is the label of $X^{1}$, whereas we address the multimodal case with $X^{1}$ and $X^{2}$ being two data modalities, demonstrating its effectiveness in both attainable and unattainable MNI scenarios.
@@ 3 Experiments
We conduct a simulation study and two real-world multimodal experiments to evaluate the efficacy of our proposed DisentangledSSL. We compare DisentangledSSL against a disentangled variational autoencoder baseline, DMVAE (Lee & Pavlovic, 2021), as well as various multimodal contrastive learning methods, including CLIP (Radford et al., 2021), which aligns different modalities to learn the shared representations, and FOCAL (Liu et al., 2024b), FactorCL (Liang et al., 2024) and SimMMDG (Dong et al., 2023), which capture both shared and modality-specific information. Additionally, we evaluate a joint optimization variant of DisentangledSSL, JointOpt (Pan et al., 2021), to highlight the advantages of our step-by-step approach.
@@ 3.1 Simulation Study
Results. We assess the performance of the learned shared and modality-specific representations for different values of $\beta$ and $\lambda$, as shown in Figure 4. For comparison, we also evaluate JointOpt, DMVAE, FOCAL, and SimMMDG across different hyperparameter settings. Specifically, for JointOpt, we vary both $a$ and $\lambda$, where $a$ controls the joint mutual information terms $I(\hat{Z}_{s}^{1},\hat{Z}_{c}^{2};X^{1})$ and $I(\hat{Z}_{s}^{2},\hat{Z}_{c}^{1};X^{2})$, and $\lambda$ adjusts the mutual information term between representations, similar to DisentangledSSL.
@@ 3.2 MultiBench
Figure 5: Results of different representations learned by DisentangledSSL.
Results. We evaluate the linear probing accuracy of representations learned by each pretrained model, as shown in Table 1. FactorCL-emb refers to the embeddings of FactorCL before projection heads, while FactorCL-proj uses the concatenation of all projection head outputs. All models use representations (or concatenation of them, if applicable) with the same dimensionality. We present the best results among three representations–shared, modality-specific, and the concatenation of both–for all methods in Table 1, with results for each representation in Appendix H.2. DisentangledSSL consistently outperforms baselines across datasets, demonstrating its ability to capture valuable information for downstream tasks. Further, we show the performance of each representation learned through DisentangledSSL in Figure 5. Combining shared and specific representations of DisentangledSSL improves performance in most cases, showing that both contribute to label prediction, while in MOSI and MUSTARD, only shared or specific information contributes significantly. For example, for MUSTARD, specific representations capture nuances of sarcasm, such as sardonic expressions or ironic tone, which are crucial for sarcasm prediction.
@@ 4 Related Work
Multimodal disentanglement in self-supervised learning. Recent studies have explored disentangled representation learning in multimodal self-supervised contexts. Liang et al. (2024) factorizes and optimizes mutual information bounds to capture the union of shared and unique features, while not penalizing for redundancy between latents. Zhang et al. (2024) learns disentangled representations of cell state for multimodal data in biological contexts. Liu et al. (2024b) and Li et al. (2024) use mutual information optimization with orthogonal/MMD regularizations for disentangling multimodal time-series signals and T-cell receptor segments, respectively. Li et al. (2022) and Dong et al. (2023) address the separation of shared and specific information, either among views of images or in the domain generalization setting. However, these methods lack a general framework for understanding information content, particularly when data modalities are deeply entangled in real-world applications.
@@ 5 Conclusions
6 Reproducibility Statement
For the theoretical results presented in the paper, we provide complete proofs in Appendix D. For the experimental results, we provide detailed explanations of the algorithm implementations and experimental setup in Section 3, Appendix G, and Appendix H. For the datasets used in the experiments, we utilize publicly available datasets and follow the data processing procedures in prior works, with additional details elaborated in Section 3 and Appendix H. The code can be found in https://github.com/uhlerlab/DisentangledSSL.
@@ Acknowledgement
XZ, ST and CU were partially supported by the Eric and Wendy Schmidt Center at the Broad Institute, NCCIH/NIH (1DP2AT012345), ONR (N00014-22-1-2116) and DOE (DE-SC0023187). SG and SJ acknowledge the support of the NSF Award CCF-2112665 (TILOS AI Institute), an Alexander von Humboldt fellowship, and Office of Naval Research grant N00014-20-1-2023 (MURI ML-SCOPE). CW and TJ acknowledge support from the Machine Learning for Pharmaceutical Discovery and Synthesis (MLPDS) consortium, the DTRA Discovery of Medical Countermeasures Against New and Emerging (DOMANE) threats program, and the NSF Expeditions grant (award 1918839) Understanding the World Through Code.
@@ Appendix B Background on Information Bottleneck
$\mathcal{L}=I(Z^{1};X^{1})-\beta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{I(Z^{1};X^{2})}}$ (6)
@@ Appendix H Experimental Details and Additional Results
For SimMMDG, we tune the hyperparameters $a$ and $\lambda$, where $a$ controls the weight of the cross-modal translation loss and $\lambda$ adjusts the distance loss between shared and specific representations. We use the same set of $a$ and $\lambda$ as that in JointOpt.
@@
In addition, we evaluated our method on synthetic data with varying noise levels (i.e. variance of the Gaussian noise and rate of random dropout). As shown in Figure 11, as the noise level decreases, the learned shared representation $\hat{Z}^{c}$ contains less modality-specific information under the same $\beta$, and the test accuracy on $Y_{1}$ and $Y_{2}$ decrease to close to 0.5 even with small $\beta$. Meanwhile, the frontier of test accuracy on $Y_{c}$ versus $Y_{1}$ shrinks towards the top left corner, indicating a decreasing expressivity-redundancy tradeoff and easier disentanglement. These trends align well with our theoretical analysis.
(a) Performance of $\hat{Z}_{c}$ on $Y_{1}$.
(b) Test accuracy on $Y_{c}$ versus $Y_{1}$.
(c) Performance of $\hat{Z}_{c}$ on $Y_{2}$.
(d) Test accuracy on $Y_{c}$ versus $Y_{2}$
Figure 11: Performance of shared representation $\hat{Z}_{c}$ learned with different noise level. As the noise level decreases, $\hat{Z}^{c}$ contains less modality-specific information, and the frontier of test accuracy on $Y_{c}$ versus $Y_{1}$ shrinks toward the top left, reflecting reduced expressivity-redundancy tradeoff.
@@
Dataset description. We utilize the real-world multimodal benchmark from MultiBench (Liang et al., 2021). We provide a brief description of each dataset as below.
  • MIMIC is a large-scale dataset in healthcare, which contains over 40,000 ICU patient records. The two modalities are the time series patient records measured by hours and tabular static data (age, gender, etc.). The binary task label is on whether the patient fits any ICD-9 code in group 7.
  • MOSEI is a sentence-level multimodal sentiment and emotion benchmark with 23,000 monologue videos. The two modalities are the vision and text modalities, and the label is on whether the sentiment is positive or negative corresponding to each video.
  • MOSI is a similar multimodal sentiment analysis dataset based on 2,199 YouTube video clips, with vision and text modalities and the sentiment label.
  • UR-FUNNY is a humor detection dataset in human speech, which consists of samples from TED talk videos, with vision and text modalities and a binary label on whether there is humor or not.
  • MUSTARD is a corpus of 690 videos from popular TV shows for sarcasm detection. The vision and text modalities are utilized to classify a binary label of sarcasm.
@@
Ablation study. We conduct an ablation study of DisentangledSSL on the MultiBench datasets to examine the impact of the hyperparameters $\beta$ and $\lambda$ in Table 5. As $\beta$ increases, the shared representations preserve only the most closely shared information and discard other features. Therefore, on datasets where the loosely shared information contributes to the label (eg., UR-FUNNY, MOSI, MIMIC), the prediction accuracy decreases with increasing $\beta$. For the MUSTARD dataset where there exist contradictions between language and video (Liang et al., 2024), the accuracy of the shared representations increases with very high $\beta$ since the contradicted parts are discarded and only clear signals remain. Similarly, as $\lambda$ increases, the independence constraint between the shared and specific representations is emphasized more. Thus, the modality-specific representations contains less information, potentially related to the label, making their prediction accuracy decrease.
Table 5: Prediction accuracy (%) of the representations learned by DisentangledSSL with different $\beta$ and $\lambda$ and standard deviations over 3 random seeds.
Dataset $\beta$ $\lambda$ Shared Specific Both
0.01 0.001 64.08 (1.48) 62.04 (0.66) 64.43 (1.14)
1.0 0.0 64.14 (1.53) 61.81 (0.08) 64.24 ( 0.96)
UR-FUNNY 1.0 0.001 64.14 (1.53) 60.27 (1.28) 64.24 (1.54)
1.0 0.1 64.14 (1.53) 60.46 (0.32) 64.46 (0.74)
100.0 0.001 62.70 (0.70) 58.76 (0.35) 61.72 ( 0.83)
0.0001 0.001 65.16 (0.60) 51.70 (0.77) 64.97 (0.34)
0.01 0.0 65.16 (0.81) 51.36 (1.37) 64.97 (0.50)
MOSI 0.01 0.001 65.16 (0.81) 51.70 (0.72) 65.11 (0.80)
0.01 0.1 65.16 (0.81) 50.53 (1.31) 64.72 (0.83)
1.0 0.001 63.31 (0.70) 50.87 (1.15) 63.95 (0.77)
0.01 0.001 62.73 (0.26) 66.05 (0.39) 66.25 (0.36)
1.0 0.0 63.16 (0.48) 65.74 (0.20) 66.26 (0.35)
MIMIC 1.0 0.001 63.16 (0.48) 65.73 (0.09) 66.44 (0.31)
1.0 0.1 63.16 (0.48) 65.37 (0.48) 66.38 (0.19)
100.0 0.001 61.94 (0.29) 62.82 (0.85) 62.70 (0.63)
0.01 0.001 76.68 (0.21) 75.71 (0.43) 77.00 (0.14)
1.0 0.0 76.94 (0.22) 75.97 (0.52) 77.54 (0.19)
MOSEI 1.0 0.001 76.94 (0.22) 75.99 (0.60) 77.45 (0.06)
1.0 0.1 76.94 (0.22) 75.83 (0.39) 77.56 (0.11)
100.0 0.001 77.20 (0.16) 75.35 (0.72) 77.26 (0.27)
0.01 0.001 56.28 (1.90) 58.70 (1.57) 55.80 (2.05)
1.0 0.0 54.11 (1.51) 60.14 (0.59) 56.28 (4.16)
MUSTARD 1.0 0.001 54.11 (1.51) 61.60 (2.61) 56.52 (2.18)
1.0 0.1 54.11 (1.51) 61.35 (2.08) 57.49 (1.49)
100.0 0.001 57.25 (3.13) 61.84 (4.78) 57.97 (2.05)